EXPERIMENT Provided Materials and Equipment: General glassware found in the lab or in your drawer/locker 60 mL of 0.1 M HCl(aFor the acid: Pour 60 mL of HCL in a buret using a funnel. Prepare 25 mL of each solution in a separate beaker (water, eye saHow Buffers Work To discuss how a buffer works, we will consider the equilibrium of a weak acid. HA(aq) + H2O(I) = A (aq) + HBuffer Capacity Effectiveness of a buffer or buffer capacity refers to the amount of acid or base that a buffer can neutralizmoles of acid needed to decrease the pH by 1unit Acid capacity of a buffer (mol L-2) = hit (16-3) volume of buffer usedin Lit1. Show the calculations you used to transform your data into a meaningful set of numbers.Acid HCl 60mL Acid Final volume Initial volume PH Acid Initial Volume PH after Final volume in heaker before in beaker 24 mL
Answer #1

In the given experiment, the aim is mostly to study and compare the differences in the acid and base buffering capacity of water, eye saline, listerine, ice tea.

Using the formula given for acid capacity of buffer and base capacity of buffer, let's first calculate the same for each of them.

Acid capacity of buffer Base Capacity of buffer
Water

24 * 0.1/ 6.36 * 25 = 0.015

10.1 * 0.1 / 6.32 * 25 = 0.0064

Eye Saline

8 * 0.1/ 6.04 * 25 = 0.0053 10.5 * 0.1 / 1.28 * 25 = 0.033
Listerine

10.3 * 0.1 / 2.35 * 25 = 0.017

18.1 * 0.1 / 9.24 * 25 = 0.0078
Ice tea

6.2 * 0.1 / 1.55 * 25 = 0.016

14.7 * 0.1 / 8.45 * 25 = 0.0069

Explaining the above calculation,

For acid capacity of water,

Volume of acid added is 24mL; 6.36 is the change in pH units; 25mL is the volume of buffer taken.

Thus, moles of acid needed to change pH by 1 unit is - (24mL* 1litre * 0.1moles/ litre ) / 1000mL * 6.36 pH units = 0.00037moles

volume of buffer taken in litre = 25/1000 = 0.025 litre

Acid capacity of buffer = 0.00037 / 0.025 = 0.015 moles / litre

Similar calculations are done for the rest too.

Thus, water, listerine and ice tea have better acid capacity of buffer. While eye saline has better base buffer capacity.